Kari is 6 years older than Adam. In 9 years, 1/2 of Adams age will equal 1/3 Kari's age. What are the ages of Adam and Kari in 2 years?
Answer by stanbon(75887)

Kari is 6 years older than Adam. In 9 years, 1/2 of Adams age will equal 1/3 Kari's age. What are the ages of Adam and Kari in 2 years?
-----
K = A+6
(1/2)A = (1/3)K
----------
Substitute for "K" and solve for "A"::
A/2 = (A+6)/3
--------
3A = 2A+12
Adams's age = 12 yrs
----
Kari's age = A+6 = 18 yrs
-----------------
Cheers,
Stan H.
-----------------------